OpenStudy (anonymous):

How would I solve the following problem. a^+b^=c^ A= 3 b= square root 5

OpenStudy (zzr0ck3r):

what does a^ mean?

OpenStudy (anonymous):

A squared

OpenStudy (anonymous):

So the equation is A^2+B^2=C^2 where A =3 and b= SQ root of 5. so replace those in the equation... 3+SqRoot 5= C^2